OpenStudy (anonymous):

how to find csc(x)=sec(x) for x from 0,2pi

OpenStudy (anonymous):

cosx=sinx <-cosx=sin(x+pi/2) sin(x+pi/2)=sinx .: x+pi/2=x ??

OpenStudy (anonymous):

x-x=-pi/2 0=-pi/2, haha.. lol nice joke^^

OpenStudy (anonymous):

i think sin(x) = cos(x) if x=pi/4 and x=5pi/4

OpenStudy (anonymous):

ye but why cant you do what i did

OpenStudy (anonymous):

U have Tan x = 1 so in the range 0 to 2pi you have x=?

OpenStudy (anonymous):

yeah why cant do my way though ??

OpenStudy (anonymous):

i mean whats wrong with what i did

OpenStudy (anonymous):

you are missing solutions, and you have made a mistake in the identity

OpenStudy (unklerhaukus):

\[\cos(x)=\sin(x)\] \[\sin(x-\pi/2)=\sin(x)\] the arc sine is not defined for both sides at the same time

OpenStudy (anonymous):

should be sin( pi/2 - x) = sin(x)

OpenStudy (anonymous):

so was the problem when i subbed sin(pi/2+x)=sinx

OpenStudy (anonymous):

i thought sin(pi/2+x)=sinx as well

OpenStudy (anonymous):

cosx*

OpenStudy (anonymous):

ah yes. i think the problem is Arcsin

OpenStudy (anonymous):

Hello - so you've got sin x = cos x . Where are the angles of such S-Y-M-M-E-T-R-Y ?

OpenStudy (anonymous):

?? so was my substitution of sin(pi/2+x) invalid

OpenStudy (unklerhaukus):

taking the arcsin of both sides was the invalid step